# Bidirectional Consistency Models

### v2

@@ Appendix B Derivation of the Network Parameterization
In this section, we provide more details about our network parameterization design. To start with, recall that in Song et al. [17], they parameterize the consistency model using skip connections as
@@
to hold at some very small noise scale $\varepsilon\approx 0$ so $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{0},\varepsilon)={\mathbf{x}}_{0}$ [^4] While the parameterization written in the original paper of Song et al. [17] did not explicitly include $c_{\text{in}}(t)$, we find it is actually included in its official implementation at https://github.com/openai/consistency_models_cifar10/blob/main/jcm/models/utils.py#L189 in the form of Equations 24 and 25.. Since we expect the output is a noise image of target noise scale $u$, we expand the parameterization of $c_{\text{skip}},c_{\text{out}}$ and $c_{\text{in}}$ to make them related to $u$, as
@@
Our derivation of network parameterization shares the same group of principles in EDM [15]. Specifically, we first require the input to the network $F_{\bm{\theta}}$ to have unit variance. Following Eq. (114) $\sim$ (117) in EDM paper [15], we have
@@
Then, as we discussed in the main text, we expect the model to achieve consistency in Equation 12 along the entire trajectory. According to Lemma 1 in Song et al. [17], we have
@@
in which we follow Song et al. [17] to estimate the expectation with ${\mathbf{x}}$ in $(i)$. When $u$ and $t$ are close, we can use the Euler solver to estimate ${\mathbf{x}}_{u}$, i.e,
@@
Therefore, when $u$ and $t$ are close, and base on Song et al. [17]’s approximation in $(i)$, we can rewrite the consistency defined in Equation 12 as
@@
For simplicity, we set $d(\cdot,\cdot)$ to $L_{1}$ norm in Equation 36. Note that, in practice, for $D$-dimensional data, we follow Song and Dhariwal [18] to use Pseudo-Huber loss $d({\mathbf{a}},{\mathbf{b}})=\sqrt{||{\mathbf{a}}-{\mathbf{b}}||^{2}+0.00054^{2}D}-0.00054\sqrt{D}$, which can be well approximated by $L_{1}$ norm.
@@
Following Karras et al. [15], we require the effective training target to have unit variance, i.e.,
@@
which is a hyperbolic function of $c_{\text{skip}}(t,u)$. Following Karras et al. [15], we select $c_{\text{skip}}(t,u)$ to minimize $|c_{\text{out}}(t,u)|$ so that the errors of $F_{\bm{\theta}}$ are amplified as little as possible, as
@@
On the side of CMs, setting $u=\varepsilon$ will arrive at exactly the same form of $c_{\text{in}}(t,u)$ and $c_{\text{out}}(t,u)$ in Equation 25. While $c_{\text{skip}}(t,u)$ does not degenerate exactly to the form in Equation 25 when taking $u=\varepsilon$, this inconsistency is negligible when $\varepsilon\approx 0$, as
@@
Therefore, we conclude that our parameterization is compatible with CM’s parameterization, so with the same CT target of Equation 11, any CT techniques [17, 18] should directly apply to our model and it should inherit all properties from CMs just by setting $u=\varepsilon$, which is a clear advantage compared with models that adopt completely different parameterizations (e.g., CTM [19]).
Appendix C Comparison of CT, CTM and BCT
@@
Appendix D Additional Ablation and Conclusions
Figure 8: Tracking the loss with different objection functions. We include the loss curve of iCT [18] for reference. We can see that the model with BCT loss defined in Equation 15 converges well. Conversely, the model applying Equation 13 instead of Equation 14 for the soft constraint has a much higher loss at the end of the optimization. While the one without CT loss totally diverges.
D.1 Comparison between Loss defined with Equation 13 and Equation 14
(a) original images
(b) Inversion w. NFE=1
(c) Inversion w. NFE=4
(d) Inversion w. NFE=12
We present the loss curve of both choices in Figure 8, where we can see that the model trained with Equation 13 features a much higher loss in the end. This echoes its failure in the inversion process: as shown in Figure 9, the model trained with Equation 13 fails to provide an accurate inversion. This is because Equation 13 contains two trajectories, starting from ${\mathbf{x}}_{u}$ and ${\mathbf{x}}_{t}$. While both of them are along the SDE trajectories starting from the same ${\mathbf{x}}_{0}$, they do not necessarily reside on the same PF ODE trajectory; in fact, the probability measure of them being on the same PF ODE trajectory is 00. On the contrary, Equation 14 bypasses this issue since it only involves trajectories with the same starting point ${\mathbf{x}}_{t}$.
D.2 Ablation of CT loss
We note that the soft constraint defined in Equation 14 can, in principle, cover the entire trajectory, it should also be able to learn the mapping from any time step $t$ to 00, which is the aim of CT loss. However, we find it crucial to include CT loss in our objective since the probability of the endpoint of the trajectory being selected is very small, as we will see in Section D.3. We provide the loss curve trained without CT loss term in Figure 8, where we can see the training fails to converge. For further verification, we also visualize the images generated by the model trained with full BCM loss and the model trained without CT loss term after 200k iterations in Figure 10. We can clearly see that the model without CT loss cannot deliver meaningful outcomes.
D.3 Coverage of Trajectory during Training
This also points out some future directions to improve BCM. For example, we can design a better sampling strategy during training to ensure a better coverage of the entire trajectory, similar to what have been done for consistency training by Song and Dhariwal [18].
Appendix E Understanding the Learned Noise (“Embedding”) Space
@@
We randomly select 500 CIFAR-10 images, and randomly split them into 10 groups. We then invert the images to their corresponding noise by Algorithm 5. We inject the same initial noise during inversion for images in the same group. Figure 12 visualize the t-SNE results [57] of the inversion outcomes. In Figure 12, images injected with the same initial noise are shown in the same color; while in Figure 12(b), we color the points according to their class label (i.e., airplane, bird, cat, …).
Interestingly, we can see that images inverted with the same initial noise are clustered together. We, therefore, conjecture that each initial noise corresponds to a submanifold in the final “embedding” space. The union of all these submanifolds constitutes the final “embedding” space, which is the typical set of $\mathcal{N}(0,T^{2}{\bm{I}})$, closed to a hypersphere. This explains why applying the same initial noise is suboptimal in interpolation, as discussed in Section A.3.1.
Appendix F Additional Results
Figure 13: Classification accuracy after BCM’s purification under different attack strengths. Different color represents different purification (defense) strengths, as we discussed in Section 4.3.4.
(a) Adversarial image with $\varepsilon_{\infty}=32/255$
(b) Adversarial image with $\varepsilon_{\infty}=8/255$
(c) Adversarial image with $\varepsilon_{\infty}=32/255$, purified by BCM with $t=80$
(d) Adversarial image with $\varepsilon_{\infty}=8/255$, purified by BCM with $t=80$
(e) Adversarial image with $\varepsilon_{\infty}=32/255$, purified by BCM with $t=1.0$
(f) Adversarial image with $\varepsilon_{\infty}=8/255$, purified by BCM with $t=1.0$
Figure 14: Examples of adversarial images and purified ones with our BCM.
@@ Appendix F Additional Results
Figure 16: Reconstructed images and their residual with unconditional BCM. We include EDM’s results in (e) and (f) for comparison.
Figure 17: Interpolation between two given images (injecting different initial noise in inversion).
Figure 18: Interpolation between two given images (injecting the same initial noise in inversion).
@@
(b) Inpainted images with BCM, where the unmasked region remains unchanged from the original (NFE=4, test set FID 12.32).
@@
Figure 19: Inpainting with BCM.
@@
Figure 20: Inpainting with CM (concretely, iCT).
(a) JPEG-compressed images.
(b) BCM-restored images.
Figure 21: Restoration of uncurated JPEG-compressed Images.
(a) INR-compressed images.
(b) BCM-restored images.
Figure 22: Restoration of uncurated INR-compressed Images.
Appendix G More Samples
(a) 1-step generation (FID=3.10)
(b) 2-step generation (FID=2.39)
(c) 3-step generation (FID=2.50)
(d) 4-step generation (FID=2.29)
Figure 23: Uncurated samples generated by BCM.
(a) 1-step generation (FID=2.68)
(b) 2-step generation (FID=2.44)
(c) 3-step generation (FID=2.28)
(d) 4-step generation (FID=2.20)
Figure 24: Uncurated samples generated by BCM-conditional. Each line corresponds to one class.
@@ Appendix G More Samples
Figure 25: Uncurated samples generated by BCM-deep.

### v4

@@ Appendix D Derivation of the Network Parameterization
In this section, we provide more details about our network parameterization design. To start with, recall that in Song et al. (2023), they parameterize the consistency model using skip connections as
@@
to hold at some very small noise scale $\varepsilon\approx 0$ so $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{0},\varepsilon)={\mathbf{x}}_{0}$ [^4] While the parameterization written in the original paper of Song et al. (2023) did not explicitly include $c_{\text{in}}(t)$, we find it is actually included in its official implementation at https://github.com/openai/consistency_models_cifar10/blob/main/jcm/models/utils.py#L189 in the form of Equations 17 and 18.. Since we expect the output is a noise image of target noise scale $u$, we expand the parameterization of $c_{\text{skip}},c_{\text{out}}$ and $c_{\text{in}}$ to make them related to $u$, as
@@
Our derivation of network parameterization shares the same group of principles in EDM Karras et al. (2022). Specifically, we first require the input to the network $F_{\bm{\theta}}$ to have unit variance. Following Eq. (114) $\sim$ (117) in EDM paper (Karras et al., 2022), we have
@@
Then, as we discussed in the main text, we expect the model to achieve consistency in Equation 7 along the entire trajectory. According to Lemma 1 in Song et al. (2023), we have
@@
in which we follow Song et al. (2023) to estimate the expectation with ${\mathbf{x}}$ in $(i)$. When $u$ and $t$ are close, we can use the Euler solver to estimate ${\mathbf{x}}_{u}$, i.e,
@@
Therefore, when $u$ and $t$ are close, and base on Song et al. (2023)’s approximation in $(i)$, we can rewrite the consistency defined in Equation 7 as
@@
For simplicity, we set $d(\cdot,\cdot)$ to $L_{1}$ norm in Equation 29. Note that, in practice, for $D$-dimensional data, we follow Song & Dhariwal (2024) to use Pseudo-Huber loss $d({\mathbf{a}},{\mathbf{b}})=\sqrt{||{\mathbf{a}}-{\mathbf{b}}||^{2}+0.00054^{2}D}-0.00054\sqrt{D}$, which can be well approximated by $L_{1}$ norm.
@@
Following Karras et al. (2022), we require the effective training target to have unit variance, i.e.,
@@
which is a hyperbolic function of $c_{\text{skip}}(t,u)$. Following Karras et al. (2022), we select $c_{\text{skip}}(t,u)$ to minimize $|c_{\text{out}}(t,u)|$ so that the errors of $F_{\bm{\theta}}$ are amplified as little as possible, as
@@
On the side of CMs, setting $u=\varepsilon$ will arrive at exactly the same form of $c_{\text{in}}(t,u)$ and $c_{\text{out}}(t,u)$ in Equation 18. While $c_{\text{skip}}(t,u)$ does not degenerate exactly to the form in Equation 18 when taking $u=\varepsilon$ and $t>\varepsilon$, this inconsistency is negligible when $\varepsilon\approx 0$, as
@@
Therefore, we conclude that our parameterization is compatible with CM’s parameterization, so with the same CT target of Equation 6, any CT techniques (Song et al., 2023; Song & Dhariwal, 2024) should directly apply to our model and it should inherit all properties from CMs just by setting $u=\varepsilon$, which is a clear advantage compared with models that adopt completely different parameterizations (e.g., CTM Kim et al. (2024)).
Appendix E Comparison of CT, CTM and BCT
@@
Appendix F Understanding the Learned Noise (“Embedding”) Space
@@
We randomly select 500 CIFAR-10 images, and randomly split them into 10 groups. We then invert the images to their corresponding noise by Algorithm 5. We inject the same initial noise during inversion for images in the same group. Figure 15 visualize the t-SNE results Van der Maaten & Hinton (2008) of the inversion outcomes. In Figure 15(a), images injected with the same initial noise are shown in the same color; while in Figure 15(b), we color the points according to their class label (i.e., airplane, bird, cat, …).
Interestingly, we can see that images inverted with the same initial noise are clustered together. We, therefore, conjecture that each initial noise corresponds to a submanifold in the final “embedding” space. The union of all these submanifolds constitutes the final “embedding” space, which is the typical set of $\mathcal{N}(0,T^{2}{\bm{I}})$, closed to a hypersphere. This explains why applying the same initial noise is suboptimal in interpolation, as discussed in Section C.3.1.
Appendix G Additional Visualizations
@@
Figure 17: Reconstructed images and their residual with unconditional BCM on CIFAR-10. We include EDM’s results in (e) and (f) for comparison.
Figure 18: Interpolation between two real CIFAR-10 images (injecting different initial noise in inversion).
Figure 19: Interpolation between two real CIFAR-10 images (injecting the same initial noise in inversion).
Figure 20: Interpolation between two real FFHQ $64\times 64$ images (injecting different initial noise in inversion).
Figure 21: Interpolation between two real LSUN Bedroom $256\times 256$ images (injecting different initial noise in inversion).
@@
(b) Inpainted images with BCM, where the unmasked region remains unchanged from the original (NFE=4, test set FID 12.32)
@@
Figure 22: Inpainting with BCM on CIFAR-10.
@@
Figure 23: Inpainting with CM (concretely, iCT) on CIFAR-10.
Figure 24: Inpainting on FFHQ $64\times 64$.
Figure 25: Inpainting on LSUN Bedroom $256\times 256$.
Appendix H More Generation Samples
@@
Figure 26: Uncurated CIFAR-10 samples generated by BCM-deep.
(a) 1-step generation (FID=3.14)
(b) 2-step generation (FID=2.45)
(c) 3-step generation (FID=2.61)
(d) 4-step generation (FID=2.35)
Figure 27: Uncurated ImageNet-64 samples generated by BCM-deep. Each line corresponds to one randomly selected class.